Al is now 3 times as old as Jay but 5 years ago, he was 5 times as old as Jay. How old is Al now?
Answer by htmentor(1343)

Let A = Al's current age
So Jay's age now = J = A/3
5 years ago, Al was 5 times as old:
A-5 = 5(J-5)
Simplify and substitute the value for J from the 1st equation:
A-5 = 5(A/3-5)
A-5 = (5/3)A - 25
(2/3)A = 20
A = 30
So Al's current age is 30